WHICH equation describes the relationship between X and Y as in the table below x y
3 5
4 9
5 13
6 17
a. y=4x-7,b.y=x+12,c.y=2x-5,d.y=3x-5
can you give me the answer and explane it for me please!!!!!!!!!!!!
Answer #1
pizza
answered by: Matt (Newton's laws)
Answer #2
i would plug in the x's and y's in each equation and see which one works...example; x=3 and y=5
a. y=4x-7 5=4(3)-7 5=12-7 =5=5 so you know that these x and y values work for this equation
b. y=x+12 5=3+12 5=15 this doesn't work so you know that those x and y values are not correct because they don't work in every equation so you have to try the next set of x's and y's: 4 and 9
hope this makes sense. it takes a while but you get the right answer
answered by: rose - spydag
Answer #3
im know whre you live
answered by: kdwk